# FreeStyle: Free Lunch for Text-guided Style Transfer using Diffusion Models

### v2

@@ 4.3 Ablation Study
Effect of hyperparameters $b$ and $s$. We present the results of ablation experiments conducted on hyperparameters $b$ and $s$, as in Figure 6. In FreeStyle, the intensities of content and style information are adjusted by the two hyperparameters $b$ and $s$, respectively. We observed that when the value of $b$ is smaller, the content in the image is more severely disturbed by the style. On the other hand, a larger value of $s$ leads to smoother image textures, while a very small $s$ value creates textures resembling noise. Overall, FreeStyle shows low sensitivity to hyperparameters, demonstrating robustness.
@@
Effect of hyperparameters $s$. We conduct ablation experiments on the transfer of “origami art” style using different settings of $s$, in Figure 7. It is evident that adjusting the hyperparameter $s$ significantly affects the intensity of the style in the images. As $s$ increases, the style intensity enhances while the content information is relatively diminished. Conversely, reducing $s$ weakens the style intensity and can even lead to inaccuracies in style expression, as seen in the second row of Figure 7 where $s=0.2$.
@@
Effect of hyperparameter $\sigma$. Figure 8 illustrates the impact of the hyperparameter $\sigma$ on the style transfer effect. The observations indicate that better style transfer are achieved when $\sigma$ exceeds 850, whereas the effect gradually deteriorates as $\sigma$ becomes too small. We believe a too small $\sigma$ value makes $f_{s}$ contain excessive content information, significantly disrupting the style information.
@@
Content-Style Disentanglement. To further validate FreeStyle’s ability to disentangle content and style information, we introduced varying degrees of $\rho$ noise into the input $x_{0}$ of the content feature $f_{c}$ to reduce content information input and observed the preservation of content and style information. As shown in Figure 9, with the increase of $\rho$ and hence more noise, the content information in $f_{c}$ gradually decreases while the style feature $f_{s}$ remains unchanged. It is clearly observed that as the value of $\rho$ increases, content information progressively decreases without affecting the expression of style information. When $\rho=999$, content information almost completely disappears, yet the expression of “sketching” style lines and brushstrokes remains observable. This validates FreeStyle’s powerful capability in disentangling content and style information.
@@ 5 Conclusion
In this study, we present FreeStyle, an innovative text-guided style transfer method that utilizes pre-trained large text-guided diffusion models. Diverging from previous approaches, FreeStyle accomplishes style transfer without the need for additional optimization or reference style images. The framework, comprising a dual-stream encoder and a single-stream decoder, seamlessly adapts to specific style transfer tasks by adjusting scaling factors. Despite its simplicity, our method showcases superior performance in terms of visual quality, artistic consistency, and robust preservation of content information across diverse styles and content images. These findings significantly advance the field of training-free style transfer.
A limitation of our algorithms is that for some challenging content images or style types, frequent manual adjustment of hyperparameters is required. However, we believe that with the diffusion model becoming more powerful and improvements in model structure, this issue will gradually diminish.

### v3

@@ 4.3 Ablation Study
Effect of hyperparameters $b$ and $s$. We present the results of ablation experiments conducted on hyperparameters $b$ and $s$, as in Fig. 6. In FreeStyle, the intensities of content and style information are adjusted by the two hyperparameters $b$ and $s$, respectively. From the experimental results, when $s$ is fixed and $b$ increases, the content information of the image becomes clearer and more complete. Conversely, when $b$ is fixed and $s$ increases, the style expression of the image becomes gradually more accurate and enhanced. However, there is also a relatively inhibitory relationship between them. When $s$ is fixed and $b$ decreases, the image shows more Ghibli-style clouds and plants. When $b$ is fixed and $s$ decreases, the results exhibit more restored content outlines. To clarify this feature more clearly, we will provide further explanation in the subsequent ablation experiment on $s$.
@@
Effect of hyperparameters $s$. We conduct ablation experiments on the transfer of “origami art” style using different settings of $s$, in Fig. 7. It is evident that adjusting the hyperparameter $s$ significantly affects the intensity of the style in the images. As $s$ increases, the style intensity enhances while the content information is relatively diminished. Conversely, reducing $s$ weakens the style intensity and can even lead to inaccuracies in style expression, as seen in the second row of Fig. 7 where $s=0.2$.
@@
Effect of hyperparameter $\sigma$. Fig. 8 illustrates the impact of the hyperparameter $\sigma$ on the style transfer effect. The observations indicate that better style transfer are achieved when $\sigma$ exceeds 850, whereas the effect gradually deteriorates as $\sigma$ becomes too small. We believe a too small $\sigma$ value makes $f_{s}$ contain excessive content information, significantly disrupting the style information. However, in our experiments, we find that setting the parameter $\sigma$ to 958 and not requiring manual tuning resulted in good performance across all images.
@@
Content-Style Disentanglement. To further validate FreeStyle’s ability to disentangle content and style information, we introduced varying degrees of $\rho$ noise into the input $x_{0}$ of the content feature $f_{c}$ to reduce content information input and observed the preservation of content and style information. As shown in Fig. 9, with the increase of $\rho$ and hence more noise, the content information in $f_{c}$ gradually decreases while the style feature $f_{s}$ remains unchanged. It is clearly observed that as the value of $\rho$ increases, content information progressively decreases without affecting the expression of style information. When $\rho=999$, content information almost completely disappears, yet the expression of “sketching” style lines and brushstrokes remains observable. This validates FreeStyle’s powerful capability in disentangling content and style information.
@@
Figure 10: The ablation study evaluates the decoupling ability of U-Net.
Figure 11: In FreeStyle, when replacing the content image with a posture image and using a person’s name as the prompt, it is clearly observable that the characters in the generated images maintain the same pose.
Ablation of U-Net Decoupling To further verify that the U-Net structure has the ability to decouple content and style, we conducte the ablation study shown in Fig. 10. Specifically, we conducte experiments by controlling variables as follows: (a) the backbone network input features are content features, and the skip connection features are replaced with null features, (b) the backbone network input features are replaced with null features, and the skip connection features are style features, (c) the backbone network input features are style features, and the skip connection features are content features, (d) the backbone network input features are content features, and the skip connection features are style features. From the experiments in Fig. 10 (a), (b), and (d), we can easily demonstrate that the model can control the output of image content and style information through the backbone features and skip connection features, respectively. In Fig. 10 (a), when the skip connection input is null, the model output still maintains the image content information unchanged. In Fig. 10 (b), when the backbone feature input is null, the model still preserves the style in the skip connection features. In Fig. 10 (c) and (d), we swapped the backbone features and skip connection features and observed the model’s generated results. We found that when content feature is used as the skip connection feature and style feature is used as the backbone feature input to the network, the model fails to generate effective results. This further verifies that the backbone features and skip connection features of U-Net correspond to the content and style information of the generated images, respectively.
4.4 Other Study
Based on the exploration and analysis of the U-Net structure in this paper, we believe that the backbone network’s ability to suppress high-frequency information and the predominance of high-frequency information in skip connection features can be used to achieve a variety of other interesting effects. Fig. 11 illustrates how we used a posture image to achieve uniform pose generation. Following this, we replace the content image with a posture image as the image input, and a person’s name replaces the style input as the prompt input. We observe that each row in the figure has generated characters consistent with the prompt, and these characters maintain the same pose as in the posture image. Since we did not include any style information in the prompt, which controls the generation of style features, we observe that the generated style exhibits noticeable inconsistency and uncertainty. In contrast, the consistency of the characters’ poses in the image is due to using the pose image as the content image input, thereby preserving the overall structural information of the image. These results indicate that this method has the potential to achieve specific pose generation after fine-tuning on a particular object.
@@ 5 Conclusion
In this study, we present FreeStyle, an innovative text-guided style transfer method that utilizes pre-trained large text-guided diffusion models. Diverging from previous approaches, FreeStyle accomplishes style transfer without the need for additional optimization or reference style images. The framework, comprising a dual-stream encoder and a single-stream decoder, seamlessly adapts to specific style transfer tasks by adjusting scaling factors. Despite its simplicity, our method showcases superior performance in terms of visual quality, artistic consistency, and robust preservation of content information across diverse styles and content images. These findings significantly advance the field of training-free style transfer. Meanwhile, as a training-free approach, the unavoidable manual parameter tuning remains an area for improvement. In future work, we will address this issue to achieve a parameter-adaptive training-free style transfer method.